OpenStudy (anonymous):

An entry in the Apple Festival Poster Contest must be rectangular and have an area of 1200 square inches. Also, its length must be 20 inches longer than its width. Find the dimensions each entry must have.

geerky42 (geerky42):

You do know that A = L W, right? We are given that: A = 1200in^2 and L = 20+W So plug in given area and substitute L to 20+W 1200 = (20+W) W Is that clear?

OpenStudy (anonymous):

Yes it is, thank you for your help.